To find the optimum shape for a cross-section of a beam against bending, beam sections of a square, circle and equilateral triangle with the same cross-sectional area (A) are considered. If the beams have the same length (L), Young’s Modulus (E) and support condition, determine which shape of the beam is the best against bending
1
Expert's answer
2019-03-01T11:27:00-0500

Beams must withstand two types of loads: elastic bending (defined by the deflection "\delta") and failure to bending (defined by the stress on the top and bottom surface "\sigma"). So calculate these values for a square, circle, and triangle with the same are A. Let's begin with deflection (smaller "\delta" is better):


"\delta=\frac{FL^2}{kEI}=\frac{G}{I},"

where G is a coefficient equal for all cross-sections and I is moment of inertia of cross-section:


"\delta_{square}=\frac{G}{A^2/12},""\delta_{circle}=\frac{G}{A^2/4\pi},"

"\delta_{triangle}=\frac{G}{A^2/6\sqrt{3}}."

We see that triangles withstand elastic bending better than squares, and squares - better than circles.

Now compute the stress inside the beams that defines failure in bending, and smaller "\sigma" is better:


"\sigma=\frac{Mc}{I}=\frac{M}{Z},"

where c - the smallest distance from the axis of a cross-section to its surface, I - moment of inertia that we already acquainted with. M - moment.

"\sigma_{square}=\frac{M}{A^{3/2}/6},"

"\sigma_{circle}=\frac{M}{A^{3/2}/4\sqrt{\pi}},"

"\sigma_{triangle}=\frac{M}{A^{3/2}/12\sqrt[4]{3}},"

that is why square resists failure in bending better than circle, circle is better than triangle.

Overall result: squares are the best, but actually one needs to choose the cross-section by what load is being considered.